A large gun fires a shell with a muzzle velocity of 400m/sec. What is the renge of the gun when it is elevated at an angle of: 30? and 50?
1
Expert's answer
2020-12-14T07:26:14-0500

Solution:

Range of object projected at an angle θ


"R=( \\dfrac{V^2}{g})sin(2\t\\theta)"


V is initial velocity in "\tfrac{m}{s}"

g is the acceleration of gravity 9.8 "( \tfrac{m}{s^2})"


"R_1=( \\dfrac{400^2}{9.8}\t)sin(2*30) = 14140m"


"R_2=( \dfrac{400^2}{9.8})sin(2*50)=16080m"